# MMVR: Millimeter-wave Multi-View Radar Dataset and Benchmark for Indoor Perception

### v1

@@ Abstract
Compared with an extensive list of automotive radar datasets that support autonomous driving, indoor radar datasets are scarce at a smaller scale in the format of low-resolution radar point clouds and usually under an open-space single-room setting. In this paper, we scale up indoor radar data collection using multi-view high-resolution radar heatmap in a multi-day, multi-room, and multi-subject setting, with an emphasis on the diversity of environment and subjects. Referred to as the millimeter-wave multi-view radar (MMVR) dataset, it consists of $345$K multi-view radar frames collected from $25$ human subjects over $6$ different rooms, $446$K annotated bounding boxes/segmentation instances, and $7.59$ million annotated keypoints to support three major perception tasks of object detection, pose estimation, and instance segmentation, respectively. For each task, we report performance benchmarks under two protocols: a single subject in an open space and multiple subjects in several cluttered rooms with two data splits: random split and cross-environment split over $395$ 1-min data segments. We anticipate that MMVR facilitates indoor radar perception development for indoor vehicle (robot/humanoid) navigation, building energy management, and elderly care for better efficiency, user experience, and safety.
@@ 1 Introduction
Compared with popular automotive radar perception efforts and datasets (Caesar et al. (2020); Schumann et al. (2021); Sheeny et al. (2021); Barnes et al. (2020); Kim et al. (2020); Ouaknine et al. (2021); Wang et al. (2021); Lim et al. (2021); Rebut et al. (2021); Palffy et al. (2022); Paek et al. (2022); Bilik et al. (2019); Pandharipande et al. (2023)), indoor radar perception receives less attention but it is essential in applications such as indoor vehicle (robot/humanoid) navigation, building energy management, and elderly care under low light and emergence (smoke, fire, and dust) situations with benefits of low device cost and less privacy concerns. As a result, its development lags and the open datasets are limited in terms of size, annotation, tasks, and benchmarks.
@@ 4 Dataset
MMVR supports a range of perception tasks: object detection, pose estimation, and instance segmentation under a diverse setting of $6$ rooms and $22$ subjects and spanning over $9$ days.
@@ 6 Responsibility to Human Subjects
Our data collection for MMVR was approved by our institutional review board (IRB). We initiated a call-for-volunteers in our institution and recruited $22$ participants. At the beginning of each data session, we informed participants about our experiment, its research goal, the procedure, potential exposure to high-frequency radio frequency waves, and the use of the camera. We notified the participants that de-identified data would be made publicly accessible for research purposes. A consent form was signed by all participants.
@@ 8 Conclusion
In this paper, MMVR scales up indoor radar data collection using multi-view high-resolution heatmaps in a multi-day, multi-room, and multi-subject setting, with an emphasis on the diversity of the environment and subjects. It complies with an extensive collection of $345$K radar heatmap frames from $22$ subjects and $6$ different rooms to benchmark three mainstream perception tasks such as object detection, pose estimation, and instance segmentation with $446$K bounding boxes/segmentation instances and $7.59$ million keypoints annotated. We hope that MMVR can stimulate the development of robust radar perception models that can operate effectively in varied real-world applications in smart home systems, security, elderly care, and navigation assistance for visually impaired individuals, contributing positively to society.
@@ 9 Contributions
M. Mahbubur Rahman, P. Wang, and P. Boufounos contributed to the development of the MMVR testbed. M. Mahbubur Rahman and P. Wang coordinated and participated in data collection. P. Li and P. Wang established the annotation pipeline for labels used in object detection, pose estimation, and segmentation with P. Wang contributing to label curation. S. Kato and P. Wang worked on the pose estimation baseline with S. Kato evaluating benchmarks under various data splits and protocols and visualizing the results. R. Yataka, A. Cardace, and P. Wang contributed to the object detection and segmentation baselines with R. Yataka evaluating benchmarks under various data splits and protocols benchmarks and visualizing the results. M. Mahbubur Rahman and P. Wang worked on the initial alignment between the camera and radar frames, the camera-radar calibration, and the signal processing pipeline to generate radar heatmaps. R. Yataka contributed to the refined alignment between camera and radar frames. S. Kato, R. Yataka, and P. Wang contributed to data segmentation, sequence splits, and the organization of the final data structure. P. Wang initiated the MMVR effort for indoor perception applications at MERL and led the project from its inception. P. Wang and P. Boufounos supervised the project.
@@ C.2 Folder Structure
Fig. 13 illustrates the folder structure of our dataset MMVR. Below the root folder, we have one data folder of $\text{d}x\text{s}y$ for each data session, where $x$ and $y$ are the indices for the day and session, respectively. Under each data folder or data session, we group the data frames into separate data segment folders named using a three-digit, zero-filled convention based on the chronological order of the data segment. For instance, the first one-minute data segment corresponds to the folder $000$, while the second data segment folder is named after $001$.
@@ G.1 Evaluation Hyper-Parameters
Original radar heatmap size $H\times W$ 512$\times$256

### v2

@@ Abstract
Compared with an extensive list of automotive radar datasets that support autonomous driving, indoor radar datasets are scarce at a smaller scale in the format of low-resolution radar point clouds and usually under an open-space single-room setting. In this paper, we scale up indoor radar data collection using multi-view high-resolution radar heatmap in a multi-day, multi-room, and multi-subject setting, with an emphasis on the diversity of environment and subjects. Referred to as the millimeter-wave multi-view radar (MMVR) dataset, it consists of $345$K multi-view radar frames collected from $25$ human subjects over $6$ different rooms, $446$K annotated bounding boxes/segmentation instances, and $7.59$ million annotated keypoints to support three major perception tasks of object detection, pose estimation, and instance segmentation, respectively. For each task, we report performance benchmarks under two protocols: a single subject in an open space and multiple subjects in several cluttered rooms with two data splits: random split and cross-environment split over $395$ 1-min data segments. We anticipate that MMVR facilitates indoor radar perception development for indoor vehicle (robot/humanoid) navigation, building energy management, and elderly care for better efficiency, user experience, and safety. The MMVR dataset is available at https://doi.org/10.5281/zenodo.12611978.
: Equal contribution. : The work of M. Rahman (Univ. of Alabama, USA), S. Kato (Osaka Univ., Japan), P. Li (Brandeis Univ., USA), and A. Cardace (Univ. of Bologna, Italy) was done during their internship at MERL. : The work was done as a visiting scientist from Mitsubishi Electric Corporation, Japan. : Project Lead.
@@ 1 Introduction
Compared with popular automotive radar perception efforts and datasets Caesar et al. (2020); Schumann et al. (2021); Sheeny et al. (2021); Barnes et al. (2020); Kim et al. (2020); Ouaknine et al. (2021); Wang et al. (2021); Lim et al. (2021); Rebut et al. (2022); Palffy et al. (2022); Paek et al. (2022); Li et al. (2022); Bilik et al. (2019); Pandharipande et al. (2023); Yataka et al. (2024), indoor radar perception receives less attention but it is essential in applications such as indoor vehicle (robot/humanoid) navigation, building energy management, and elderly care under low light and emergence (smoke, fire, and dust) situations with benefits of low device cost and less privacy concerns. As a result, its development lags and the open datasets are limited in terms of size, annotation, tasks, and benchmarks.
@@ 4 Dataset
MMVR supports a range of perception tasks: object detection, pose estimation, and instance segmentation under a diverse setting of $6$ rooms and $25$ subjects and spanning over $9$ days.
@@ 5.4.3 Comparison of Detection Backbones
Table 7: Baseline benchmarks between HIBER Wu et al. (2023) and MMVR.
Room Setting Dataset Object Detection Instance Seg.
AP AP50 AP75 IoU
Open HIBER (Walk) 17.77 52.46 6.78 48.47
Foreground MMVR (P1) 24.46 66.82 11.22 67.03
5.4.4 Comparison over Different Radar Datasets
An interesting ablation study involves training the radar perception model on one radar dataset, such as MMVR, and evaluating its performance on another dataset, like HIBER, or vice versa. However, this study requires dataset-dependent 3D radar-camera coordinate calibration and 3D-to-2D projection in the camera coordinate, due to varying relative geometries between the camera and radar and the use of different cameras in different datasets. Instead, we directly compare the same tasks across different datasets under similar environmental settings: a single subject walking in an open-foreground room. This is achieved using the “walk” data split in HIBER and “P1” in our MMVR. As shown in Table 7, the performance between the two datasets is comparable. Note that we used refined bounding boxes in the 2D image plane for the HIBER evaluation as the original bounding boxes were excessively large around the subjects.
@@ 6 Responsibility to Human Subjects
Our data collection for MMVR was approved by our institutional review board (IRB). We initiated a call-for-volunteers in our institution and recruited $25$ participants. At the beginning of each data session, we informed participants about our experiment, its research goal, the procedure, potential exposure to high-frequency radio frequency waves, and the use of the camera. We notified the participants that de-identified data would be made publicly accessible for research purposes. A consent form was signed by all participants.
@@ 8 Conclusion
In this paper, MMVR scales up indoor radar data collection using multi-view high-resolution heatmaps in a multi-day, multi-room, and multi-subject setting, with an emphasis on the diversity of the environment and subjects. It complies with an extensive collection of $345$K radar heatmap frames from $25$ subjects and $6$ different rooms to benchmark three mainstream perception tasks such as object detection, pose estimation, and instance segmentation with $446$K bounding boxes/segmentation instances and $7.59$ million keypoints annotated. We hope that MMVR can stimulate the development of robust radar perception models that can operate effectively in varied real-world applications in smart home systems, security, elderly care, and navigation assistance for visually impaired individuals, contributing positively to society.
@@ 9 Contributions
M. Rahman, P. Wang, and P. Boufounos contributed to the development of the MMVR testbed. M. Rahman and P. Wang coordinated and participated in data collection. P. Li and P. Wang established the annotation pipeline for labels used in object detection, pose estimation, and segmentation with P. Wang contributing to label curation. S. Kato and P. Wang worked on the pose estimation baseline with S. Kato evaluating benchmarks under various data splits and protocols and visualizing the results. R. Yataka, A. Cardace, and P. Wang contributed to the object detection and segmentation baselines with R. Yataka evaluating benchmarks under various data splits and protocols benchmarks and visualizing the results. M. Rahman and P. Wang worked on the initial alignment between the camera and radar frames, the camera-radar calibration, and the signal processing pipeline to generate radar heatmaps. R. Yataka contributed to the refined alignment between camera and radar frames. S. Kato, R. Yataka, and P. Wang contributed to data segmentation, sequence splits, and the organization of the final data structure. P. Wang initiated the MMVR effort for indoor perception applications at MERL and led the project from its inception. P. Wang and P. Boufounos supervised the project.
@@ C.2 Folder Structure
Fig. 13 illustrates the folder structure of our dataset MMVR. Below the root folder, we have one data folder of $\text{d}x\text{s}y$ for each data session, where $x$ and $y$ are the indices for the day and session, respectively. Under each data folder or data session, we group the data frames into separate non-overlapping data segment folders named using a three-digit, zero-filled convention based on the chronological order of the data segment. For instance, the first one-minute data segment corresponds to the folder $000$, while the second data segment folder is named after $001$.